OpenStudy (anonymous):

HELP! I GIVE MEDADAL ;) write the standard equation form for a circle B with center (2 2) and radius 4 The point (1 -3) is on a circle with a center (-2 1) write the equation of circle C write the standar equation for circle A(-1 -2)

sam (.sam.):



sam (.sam.):

B\[(x-h)^2+(y-k)^2=r^2\] \[(x-2)^2+(y-2)^2=16\]

sam (.sam.):

So, for C, we've got center (-2 1)

OpenStudy (anonymous):

ok

sam (.sam.):

\[(x+2)^2+(y-1)^2=R^2\]

sam (.sam.):

So when R=25, the circle touches the point (1 -3)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

write the standar equation for circle A(-1 -2) @.Sam.

OpenStudy (anonymous):

@.Sam. could you help me write the standared equarion for circle A(-1 -2)

OpenStudy (anonymous):

@shubhamsrg could you help me write the standatd equation for cicrcle A(-1 -2)